# Exploring Prompting Methods for Mitigating Class Imbalance through Synthetic Data Generation with Large Language Models

### v2

@@ A.2 Comparative analysis of classification performance using only synthetic data
We compare the classification performance using only synthetic data for training, as shown in Table 7. For both the Travel and HELOC datasets, the classification model trained on synthetic datasets generated by our model outperform the one trained on original data in terms of F1 score and Sensitivity. This improvement in performance suggests that the data generated by our method accurately represents the distribution of the original data. Furthermore, it highlights the effectiveness of our class-balanced generation approach in improving the performance of classification models.
Table 8: Comparison of token efficiency and generation stability for ablating prompt design elements on the Sick dataset
@@
Table 9: Comparison of token efficiency and generation stability for task specification in prompt design on the Sick dataset.
@@
Appendix B Additional ablation study results
This section presents additional ablation study results on the Sick datasets. We evaluate the token efficiency and generation stability of ablated versions based on 100 iterations, each with 20 random input samples. As shown in Table 8, for the Sick dataset, due to its repetitive monotonous values, the number of generated samples is low, even with grouping. Using a unique variable mapper in such cases significantly increases the number of samples generated per attempt from 6.93 to 17.68. This improvement demonstrates that our approach effectively controls the quantity of generated data samples, ensuring the stable production of synthetic tabular data. Moreover, the ability to generate multiple data samples in a single inference significantly can lowers generation time and costs.
@@ Appendix B Additional ablation study results
Figure 5: The differences between Pearson correlation matrices of real data and synthetic data for numerical variables in the Sick dataset. More intense colors indicate larger differences, with positive differences shown in red and negative differences shown in blue.
Appendix C Analysis of feature correlation between numerical variables
@@
In the main manuscript, we visualize the feature correlation between categorical variables. Here, we analyze the Pearson correlation between numerical variables in the Sick dataset and compare the differences in feature correlation values to the original data, as shown in Fig. 5. As a result, our methodology consistently shows feature correlations similar to the original data for both minor and major classes. In contrast, other baselines either show significant deviations in feature correlation for all classes or perform well for major classes but not for minor ones. These results demonstrate that our methodology shows superior performance compared to the baselines in generating high-quality data that accurately represents both minor and major classes within the original dataset.
Appendix D Impacts of sample size on the classification performance
@@
Figure 6: Impact of sample size on the classification performance on the Income and Diabetes datasets when classifiers are trained exclusively on synthetic samples generated from our proposed method. Black dashed lines denote the F1 score using only the original data.
Figure 7: Complete generation results on an imbalanced toy dataset. Major class is shown in blue and minor class in red. Our approach via in-context learning of LLM achieves (1) clear class boundaries, (2) accurate feature correlations, (3) proper value ranges, (4) well-represented relationships between categorical (last row) and numerical variables, and (5) better diversity compared to the fine-tuned GReaT [4] model.
@@
We investigate how varying the number of synthetic samples affects classification performance. This analysis is conducted on the Income and Diabetes datasets, and the results are shown in Fig. 6. For the Income dataset, sample sizes range from 500 to 10,000, while for Diabetes, the range was 500 to 20,000. For both datasets, a consistent increase in the F1 score is observed as the number of samples grow. This result demonstrates that our method can adequately represent the original data distribution by generating a sufficient number of synthetic data samples.
Figure 8: The illustration of the input samples (blue and orange) and the generated samples (green and red) for each inference on the imbalanced toy dataset when the our proposed approach received 180 input samples. Empty spaces in the visualization indicate instances where the model failed to generate samples. To produce a total of 1000 samples, the model performed 23 inferences.
Figure 9: The illustration of the input samples (blue and orange) and the generated samples (green and red) for each inference on the imbalanced toy dataset when the ablated model without grouping received 180 input samples. Empty spaces in the visualization indicate instances where the model failed to generate samples. To produce a total of 1,000 samples, the model performed 37 inferences.
Appendix E Complete qualitative results on toy dataset
@@
In Fig. 7, we illustrate the complete qualitative results on the toy dataset. We assume a challenging multivariate data scenario in an imbalanced binary classification situation, where variables with and without correlations are mixed. Using the scikit-learn library  [22], we created a total of 12 variables. These variables have strong correlations in pairs. There are 11 numerical variables and 1 categorical variable, and the categorical variable is shown on the x-axis in the last row of Fig. 7. In the figure, the major class is shown in blue, and the minor class is shown in red.
@@
The generated results can be evaluated from four aspects. Firstly, whether each data sample is correctly generated to belong to each class, i.e., whether the class boundary is clear. Secondly, whether the shapes formed by the correlation between the variables, such as circles or moons, are well represented. Thirdly, whether the values generated fall within the acceptable range for each variable. Finally, whether the relationship between the categorical and numerical variables is well represented.
@@
Overall, CSV-style generation produces better class separation boundaries compared to sentence-style generation. Regarding class presentation, giving a single class can sometimes result in good generation outcomes, but it fails to accurately capture the correlation with the major class, resulting in shapes like squares instead of circles, and it also fails to accurately generate the categorical variable. On the other hand, when generating multi-class samples, if the sample numbers are not balanced, the generation results for the minor class can be compromised. Even when balanced, providing grouping leads to clearer class distinctions. Additionally, adding unique variable mapping significantly improves the performance, resulting in a well-distinguished categorical variable.
@@
In addition, we visualize the inputs and outputs of both our method and the ablated model without grouping, as shown in Fig. 8 and Fig. 9. For each inference, 180 class-balanced samples are input into the LLM prompt. To generate a total of 1,000 samples, our method required 23 inferences, whereas the ablated model needed 37 inferences, demonstrating the stability of our method in generating synthetic data samples. In terms of data generation quality, our method successfully generates both major and minor classes in most inferences, with the distribution of the generated samples closely mirroring that of the actual input data. In contrast, the ablated model exhibits a higher number of inferences where the generated data deviates from the input data distribution.
Appendix F Broader Impact
Our method is generally applicable to any tabular classification task with minimal pre-processing, making it accessible for various research and industrial applications. The potential positive impacts include enhanced data accessibility and privacy preservation, which can improve decision-making processes in fields like healthcare and finance. Additionally, it reduces the need for using real data, thereby protecting individual privacy. However, there are potential negative impacts, such as the risk of data misuse. Synthetic data generation could be misused for malicious purposes, such as creating fake datasets to deceive systems or individuals.
Figure 10: The illustration of the input (blue and orange) and generated (green and red) samples from our method where input samples are exclusively taken from either the left or right side of the circle.
Appendix G Limitations
If the input samples do not adequately represent the data, the LLM may fail to generate samples that follow the original data distribution, leading to degraded performance. As shown in Fig. 10, if the input samples do not fully capture the original data shape, the generated data may fail to accurately follow this shape. A limitation of our work is that our current method relies on providing a large number of data samples to ensure sufficient representation in in-context learning. Future work will focus on increasing efficiency by selectively choosing key samples that can help the LLM generate data effectively with fewer examples.
Appendix H Related work
In the field of tabular data synthesis, traditional methods have predominantly focused on sampling techniques, such as SMOTE [6], which have been foundational in addressing class imbalances in tabular data. Although useful, these methods often fall short in capturing complex inter-feature correlations [2] and typically lack the versatility needed for regression tasks or addressing multiple feature imbalances. Recently, deep-learning-based generative models have demonstrated their efficacy in generating realistic tabular data [7, 9, 15, 30, 33, 34]. For example, CTAB-GAN [33] introduced a generator that mitigates the problem caused by imbalanced discrete variables. Additionally, diffusion models have also exhibited promising results in tabular data generation. TabDDPM [18] introduced a diffusion process for both numerical and categorical features in tabular data for accurate data generation. Large language model (LLM)-based methods have emerged as a recent trend in tabular dataset generation. GReaT [4] surpassed previous methods by fine-tuning LLMs to generate tabular data formatted in natural text. However, this fine-tuning process is resource-intensive, requiring extensive training of large models for each dataset.
As an alternative, leveraging the in-context learning capabilities of LLMs presents a promising solution. LLMs have demonstrated the ability to solve various problems, learn and mimic patterns from given input data, and generate new data without additional training for parameter updates [5, 13, 17, 19, 28, 29, 32]. Despite this potential, there is limited research on what LLMs generate for synthetic tabular data and what the optimal prompts for high-quality tabular data generation would be. Previous studies [31, 25] have shown the potential of LLMs for in-context learning of tabular data generation. However, these studies have not sufficiently explored optimal prompts for synthetic tabular data generation, leading to instability in the generated samples. This instability necessitates additional validation modules and a two-stage structure.
Given that the performance of in-context learning with LLMs heavily depends on prompt design [10, 26, 28, 32], carefully designing prompts specifically for imbalanced tabular data is essential to maximize LLM capabilities in effectively generating realistic data. Recognizing the importance of optimal prompt design and the potential instability in generated data, we propose a novel method for generating high-quality synthetic tabular data through tailored prompting. The proposed method is generally applicable to various types of tabular datasets without the need for training, thereby offering a cost-effective solution in data synthesis.
Appendix I Additional experimental details
@@
This section provides the illustration of ablated versions (Appendix I.1), comprehensive information on the datasets (Appendix I.2) and comparison baselines (Appendix I.3), implementation details (Appendix I.4), and final prompt design and examples (Appendix I.5).
I.1 Ablation prompt examples
@@
Figure 11: Illustration of ablated versions.
I.2 Dataset
The detailed description of the datasets are shown in Table 10. The variable descriptions are directly sourced from dataset platforms like Kaggle or the UCI repository. In our approach, we designate the class with fewer instances as the positive class for binary classification tasks.
Table 10: Details of the datasets used to evaluate synthetic tabular data generation models.
@@ I.2 Dataset
I.3 Baseline reproducibility
In our research, we utilize implementations from the Synthetic Data Vault [21] for TVAE [30], CTGAN [30], and CopulaGAN, while for the GReaT [4] model, we employ code from its official repository. In our experiment, the GReaT model fails to accurately generating column names ‘max_glu_serum’ and ‘A1Cresult’ in the Diabetes dataset, even after training for 85 epochs that reported in the original paper. Extending the training to 105 epochs does not rectify this issue, with ‘max_glu_serum’ and ‘A1Cresult’ being correctly generated in only approximately 2.8% and 12.1% of instances, respectively. For these columns, failure cases are treated as NaN value.

### v4

@@ A.2 Comparative analysis of replacing the original dataset with synthetic data
We compare the classification performance using only synthetic data for training, as shown in Table 7. For both the Travel and HELOC datasets, the classification model trained on synthetic datasets generated by our model outperforms the one trained on original data in terms of F1 score and sensitivity. This improvement in performance suggests that the data generated by our method accurately represents the distribution of the original data. Furthermore, it highlights the effectiveness of our class-balanced generation approach in improving the performance of classification models.
A.3 Additional ablation study results
This section presents additional ablation study results on the Sick datasets. We evaluate the token efficiency and generation stability of ablated versions based on 100 inferences, each with 20 random input samples. As shown in Table 8, for the Sick dataset, due to its repetitive, monotonous values, the number of generated samples is low, even with grouping. Using a unique variable mapper in such cases significantly increases the number of samples generated per attempt from 6.93 to 17.68. This improvement demonstrates that our approach effectively controls the quantity of generated data samples, ensuring the stable production of synthetic tabular data. Moreover, the ability to generate multiple data samples in a single inference can significantly lower generation time and costs.
Table 8: Comparison of token usage and generation efficiency across ablated methods on the Sick dataset. Results are based on 100 inferences, each with 20 random input samples. Input tokens indicates the number of tokens required in the LLM prompt for a fixed number of input samples. Output samples shows the average number of synthetic samples generated per iteration. Success rate measures the ratio of inferences that generate at least one valid data sample.
@@ A.3 Additional ablation study results
Table 9: Comparison of token usage and generation efficiency for task specification ablation on the Sick dataset. Results are based on 100 inferences, each with 20 random input samples.
@@
Figure 5: Difference between Pearson correlation matrices of real and synthetic data for numerical variables in the Sick dataset. More intense colors indicate larger differences, with positive differences shown in red and negative differences shown in blue. Black indicates where the correlation is not measured since only one unique value is generated for those variables.
A.4 Analysis of feature correlation between numerical variables
@@
In the main manuscript, we visualize the feature correlation between categorical variables. Here, we analyze the Pearson correlation between numerical variables in the Sick dataset and compare the differences in feature correlation values to the original data, as shown in Fig. 5. As a result, our methodology consistently shows feature correlations similar to the original data for both minority and majority classes. In contrast, other baselines either show significant deviations in feature correlation for all classes or perform well for majority classes but not for minority ones. These results demonstrate that our methodology shows superior performance compared to the baselines in generating high-quality data that accurately represents both minority and majority classes within the original dataset.
Appendix B Additional analysis of EPIC
This section provides an in-depth analysis of the proposed method, EPIC. Appendix B.1 analyzes the effect of sample size on classification performance. Appendix B.2 discusses the significance of the observed performance improvements. Appendix B.3 provides complete qualitative results on the toy dataset. Finally, Appendix B.4 explores tabular data classification with in-context learning methods.
B.1 Impacts of sample size on the classification performance
Figure 6: Classification performance on the Income dataset when synthetic data generated by our proposed method are added to the original dataset. We experiment with varying synthetic sample sizes, comparing data sampling methods: with replacement (w/ rep.) and without replacement (w/o rep.). EPIC samples with replacement. GPT-3.5-turbo is used for the experiment.
@@ B.1 Impacts of sample size on the classification performance
Figure 7: Classification performance on the Income and Diabetes datasets using only synthetic data generated by our proposed method for training. Black dashed lines denote the F1 score using only the original data. GPT-3.5-turbo is used for our method.
Adding synthetic data to the original dataset    Data samples in prompts can be selected either with or without replacement. Our method uses sampling with replacement to generate a large and diverse dataset, as sampling without replacement limits the amount of synthetic data to the number of actual data points. While we ensure that there are no overlapping examples within each prompt to maintain diversity, each prompt is constructed using sampling with replacement.
We conduct experiments to analyze (1) how much datasets can be enlarged and (2) how this impacts classification performance by comparing sampling with and without replacement. As shown in Fig. 6, performance improves steadily in both scenarios as the volume of generated data increases. However, the improvement is constrained when sampling without replacement due to the limited number of possible samples. When sampling with replacement, as the dataset size expands, there is a noticeable improvement in the balance between sensitivity and specificity, which contributes to enhanced overall performance, including gains in balanced accuracy and F1 score. Generating up to 40K synthetic data points resulted in even better performance than the 20K synthetic data points reported in Table 1 of our main manuscript. We also observed that as the volume of generated data continues to increase, the gains in balanced accuracy and F1 score eventually plateau, indicating diminishing returns and suggesting that further data generation beyond a certain point offers limited additional benefit.
@@
Using only synthetic data    We investigate how varying the number of synthetic samples affects classification performance. This analysis is conducted on the Income and Diabetes datasets, and the results are shown in Fig. 7. For the Income dataset, sample sizes range from 500 to 10,000, while for Diabetes, the range was 500 to 20,000. For both datasets, a consistent increase in the F1 score is observed as the number of samples grows. This result demonstrates that our method can adequately represent the original data distribution by generating a sufficient number of synthetic data samples.
Figure 8: Average classification results across six datasets compared with the closest baselines.
Table 10: Performance improvement after adding synthetic data to the original dataset. Results are reorganized from Table 1 of our main manuscript.
Method Dataset Improvement from the original data (%p)
F1 score $\uparrow$ BAL ACC $\uparrow$ Sensitivity $\uparrow$ Specificity $\uparrow$
GReaT Travel 2.83 1.86 1.8 1.92
Sick -0.58 -0.39 -0.76 -0.01
HELOC -0.66 -0.25 -1.67 1.18
Income 1.05 1.06 3.41 -0.31
Diabetes -0.09 -0.09 -0.02 -0.12
Thyroid -2.92 -2.62 -5.23 0.00
TabDDPM Travel -4.92 -3.3 -6.6 0.00
Sick -2.64 -1.92 -3.81 -0.04
HELOC -0.36 -0.32 -0.38 -0.26
Income -0.05 0.05 0.42 -0.31
Diabetes -0.23 -0.24 -0.09 -0.18
Thyroid 0.16 1.18 4.31 -1.96
Ours Travel 8.53 7.23 21 -6.54
Sick 0.9 1.71 3.58 -0.17
HELOC 0.91 0.45 2.07 -1.17
Income 2.26 2.7 9.17 -3.76
Diabetes 0.07 0.07 0.04 0.09
Thyroid 0.57 1.31 4.09 -1.47
B.2 Significance of performance improvements
We analyze the significance of performance improvements achieved by EPIC. As shown in Fig. 6, while the gains in F1 score and balanced accuracy between 0 and 40K samples may seem modest, the sensitivity actually increases significantly from around 50% to 70%. This improvement balances performance across all classes, leading to a more meaningful outcome and greatly enhancing model usability. Baselines often learn biases in the training data, resulting in abnormally high specificity at the expense of sensitivity. This imbalance can inflate balanced accuracy or F1 scores, but such performance is ineffective for real-world tasks. In contrast, our method significantly enhances sensitivity with only a slight reduction in specificity. As a result, our method achieves the most balanced performance across all four metrics, as shown in Fig. 8.
Our method is the only approach among the seven baselines that consistently improves both F1 score and balanced accuracy compared to using only the original data (Table 19). Extensive experiments across six real-world datasets from diverse domains (finance, healthcare, marketing, and social science) using four classifiers, each tested five times, demonstrate the state-of-the-art performance of our method, with significant gains in the highly imbalanced Travel, Sick, Income, and Thyroid datasets (Fig. 13). While GReaT and TabDDPM are the closest baselines, they exhibit inconsistent performance. As highlighted in Table 10, they reduce original data performance in more than half of the cases (red). For example, both underperform on HELOC and Diabetes.
In stark contrast, our method consistently outperforms the original data across all datasets (blue). In the challenging Diabetes dataset, only our method improves over the original data. While specificity decreases in all methods, our method still achieves a superior balance of metrics, leading to greater practicality. These results underscore that our method offers the greatest practical utility and a meaningful performance advantage among the models tested.
Figure 9: Complete generation results on an imbalanced toy dataset with majority and minority classes. Our approach, leveraging in-context learning with LLMs, achieves (1) distinct class boundaries, (2) accurate feature correlations, (3) well-matched value ranges, (4) robust numerical-categorical relationships (last row), and (5) comprehensive data distribution coverage, with improvements over its ablated versions and the fine-tuned GReaT model [4].
B.3 Complete results on toy dataset
@@
In Fig. 9, we illustrate the complete qualitative results on the toy dataset. We assume a challenging multivariate data scenario in an imbalanced binary classification situation, where variables with and without correlations are mixed. Using the scikit-learn library  [23], we created a total of 12 variables. These variables have strong correlations in pairs. There are 11 numerical variables and one categorical variable, and the categorical variable is shown on the x-axis in the last row of Fig. 9. In the figure, the majority class is shown in blue, and the minority class is shown in red.
@@
The generated results are evaluated from four aspects: (1) assessing whether generated data samples correctly belong to their respective classes, ensuring clear class boundaries; (2) examining whether the shapes formed by variable correlations, such as circles or moon patterns, are accurately represented; (3) evaluating whether the generated values fall within an acceptable range for each variable; and (4) analyzing whether relationships between categorical and numerical variables are accurately captured.
@@
Figure 10: Illustration of the input samples (majority and minority) and the corresponding generated samples (majority and minority) for each inference of EPIC on the imbalanced toy dataset. In this example, with 180 input samples per inference, EPIC requires 23 inferences to generate 1,000 samples. Empty boxes represent cases where the model fails to generate valid samples.
Figure 11: Illustration of the input samples (majority and minority) and the corresponding generated samples (majority and minority) for each inference of the ablated version of our method on the imbalanced toy dataset. In this example, with 180 input samples per inference, the ablated method without grouping requires 37 inferences to generate 1,000 samples. Empty boxes represent cases where the model fails to generate valid samples.
@@
Overall, CSV-style generation produces better class separation boundaries compared to sentence-style generation. Regarding class presentation, giving a single class can sometimes result in good generation outcomes, but it fails to accurately capture the correlation with the majority class, resulting in shapes like squares instead of circles, and it also fails to accurately generate the categorical variable. On the other hand, when generating multi-class samples, if the sample numbers are not balanced, the generation results for the minority class can be compromised. Even when balanced, providing grouping leads to clearer class distinctions. Additionally, adding unique variable mapping significantly improves the performance, resulting in a well-distinguished categorical variable.
@@
In addition, we visualize the inputs and outputs of both our method and the ablated model without grouping, as shown in Fig. 10 and Fig. 11. For each inference, 180 class-balanced samples are input into the LLM prompt. To generate a total of 1,000 samples, our method required 23 inferences, whereas the ablated model needed 37 inferences, demonstrating the stability of our method in generating synthetic data samples. In terms of data generation quality, our method successfully generates both majority and minority classes in most inferences, with the distribution of the generated samples closely mirroring that of the actual input data. In contrast, the ablated model exhibits a higher number of inferences where the generated data deviates from the input data distribution.
B.4 Comparison of advanced tabular classification models
The ability of LLMs to generate high-quality synthetic data for imbalanced classes in our approach highlights their potential to directly address classification tasks, offering a promising avenue for future research. However, the challenge lies in designing effective prompts to fully leverage this potential. The performance of LLMs is highly dependent on prompt design, and numerous studies have shown that small variations in prompt crafting can lead to significant differences in outcomes [18, 25]. Our work shares this motivation, focusing on designing prompts that enable LLMs to effectively generate high-quality tabular data, particularly to address class imbalance. While exploring LLMs for direct classification could enhance performance, this direction is beyond the scope of our current study.
We conduct experiments using recent in-context learning-based tabular classification methods, such as TabPFN [14], which employs a pretrained transformer, and T-Table [33], which leverages LLMs. We also test TabR [12], an advanced deep-learning tabular classification model, using its official code. As detailed in Table 11, we evaluate classification performance on the Travel dataset using the original and synthetic data. For T-Table, we use the GPT-3.5-turbo model, incorporating all original data but limiting synthetic samples to 200 per input prompt due to token constraints. To further enhance performance, we employ voting across five inferences [31].
Table 11: Comparison of F1 scores using robust tabular classification models on the Travel dataset. GB refers to gradient boosting classifier.
Model Original +Ours +TabDDPM +GReaT
XGBoost 55.32 67.74 51.06 64.00
LightGBM 60.00 64.29 55.32 62.50
CatBoost 57.14 64.41 48.09 57.73
GB 60.00 70.18 58.33 59.57
TabPFN 56.00 59.26 55.32 53.06
T-Table 21.62 30.43 10.53 17.65
TabR 46.41 60.78 44.88 32.41
The results indicate that, while the newly introduced classifiers do not outperform traditional ones, adding synthetic data generated by our method to the original data consistently leads to significant performance improvements across all classifiers. In contrast, using TabDDPM and GReaT often results in decreased performance. Our method uniquely and consistently enhances the performance of different classifiers, demonstrating superior label-matching quality. These findings underscore the value of high-quality synthetic data in enhancing classifier performance across diverse models, highlighting the importance of data generation research as a distinct area from classifier development.
Tabular data generation is a critical area of research with significant implications [4, 27, 37]. This task serves two primary purposes: (1) enhancing classification performance in a model-agnostic manner through data augmentation, similar to SMOTE, as shown in Tables 1, 2, and 6, and (2) generating synthetic data to replace original data in security-critical or privacy-sensitive contexts, as demonstrated in Table 7. In fields such as healthcare, where collecting new samples or achieving balanced class labels is challenging and data may be noisy or incomplete, generating high-quality synthetic data is crucial.
Figure 12: Illustration of ablated versions of our proposed method.
In conclusion, even as more advanced classification models for tabular data are developed in the future, our proposed method could continue to play a crucial role in enhancing performance by generating high-quality synthetic data, thereby potentially making a significant impact on the tabular data classification community and related tasks.
Appendix C Additional experimental details
@@
This section provides the illustration of ablated versions (Appendix C.1), comprehensive information on the datasets (Appendix C.2) and comparison baselines (Appendix C.3), implementation details (Appendix C.4), and final prompt design and examples (Appendix C.5).
C.1 Ablation prompt examples
@@
Table 12: Dataset details used in this study.
@@
C.2 Dataset
A detailed description of the datasets is shown in Table 12. The variable descriptions are directly sourced from dataset platforms like Kaggle or the UCI repository. In our approach, we designate the class with fewer instances as the positive class for binary classification tasks. We also provide the class distribution information for the datasets used in this work, as shown in Fig. 13. Among the six datasets, the Travel, Income, Sick, and Thyroid datasets exhibit substantial class imbalances.
Figure 13: Class distribution for the real-world public tabular datasets used in this study.
C.3 Baseline reproducibility
In our research, we utilize implementations from the Synthetic Data Vault [22] for TVAE [35], CTGAN [35], and CopulaGAN, while for the GReaT [4] model, we employ code from its official repository. In our experiment, the GReaT model fails to accurately generate column names ‘max_glu_serum’ and ‘A1Cresult’ in the Diabetes dataset, even after training for 85 epochs, as reported in the original paper. Extending the training to 105 epochs does not rectify this issue, with ‘max_glu_serum’ and ‘A1Cresult’ being correctly generated in only approximately 2.8% and 12.1% of instances, respectively. For these columns, failure cases are treated as NaN values.